### 28 Feb 2024 v3

# OLMo[Uncaptioned image] : Accelerating the Science of Language Models

Abstract
Language models (LMs) have become ubiquitous in both NLP research and in commercial product offerings. As their commercial importance has surged, the most powerful models have become closed off, gated behind proprietary interfaces, with important details of their training data, architectures, and development undisclosed. Given the importance of these details in scientifically studying these models, including their biases and potential risks, we believe it is essential for the research community to have access to powerful, truly open LMs. To this end, this technical report details the first release of OLMo, a state-of-the-art, truly Open Language Model and its framework to build and study the science of language modeling. Unlike most prior efforts that have only released model weights and inference code, we release OLMo and the whole framework, including training data and training and evaluation code. We hope this release will empower and strengthen the open research community and inspire a new wave of innovation.
[Uncaptioned image] Weights https://huggingface.co/allenai/OLMo-7B
[Uncaptioned image] Code https://github.com/allenai/OLMo
[Uncaptioned image] Data https://huggingface.co/datasets/allenai/dolma
[Uncaptioned image] Evaluation https://github.com/allenai/OLMo-Eval
[Uncaptioned image] Adaptation https://github.com/allenai/open-instruct
[Uncaptioned image] W&B Logs https://wandb.ai/ai2-llm/OLMo-7B/reports/OLMo-7B--Vmlldzo2NzQyMzk5
1 Introduction
Language models have been at the center of NLP technologies for many years (Rosenfeld, 2000; Bengio et al., 2003; Mikolov et al., 2013; Peters et al., 2018; Brown et al., 2020). Recently, due to large-scale pretraining and human annotation for alignment, they have become commercially valuable (OpenAI, 2023). However, as their commercial value has increased, the largest models have become gated behind proprietary interfaces, with important details left undisclosed.
We believe that full access to open language models for the research community is critical to the scientific study of these models, their strengths and weaknesses, and their biases and risks. Accordingly, we introduce OLMo, a state-of-the-art, truly open language model and framework to build, study, and advance LMs, along with the training data, training and evaluation code, intermediate model checkpoints, and training logs.
Recent LM releases have varied in their degree of openness. For example, Mistral 8x7B provided model weights and a brief report (Jiang et al., 2024), while LLaMA came with in-depth adaptation training instructions (Touvron et al., 2023b), and Mosaic Pretrained Transformer came with many details, including the dataset distribution, though not the data itself (MosaicML NLP Team, 2023). Falcon’s pretraining data was partially released (Almazrouei et al., 2023), and the most open models—the Pythia suite (Biderman et al., 2023) and BLOOM (BigScience et al., 2022)—released training code, model checkpoints, training data and more.
With OLMo, we release the whole framework from data to training to evaluation tools: multiple training checkpoints across multiple hardware types, training logs, and exact datasets used, with a permissive license. We are not the only team to do this; recent work from LLM360 targets similar goals (Liu et al., 2023). OLMo narrows the gap from their models to state-of-the-art capabilities of models like LLaMA2. This project has benefited from lessons learned from all of these previous efforts with their varying degrees of openness, and we believe that a large, diverse population of open models is the best hope for scientific progress on understanding language models and engineering progress on improving their utility.
The OLMo framework encompasses the tools and resources required for building and researching language models. For training and modeling, it includes full model weights, training code, training logs, ablations, training metrics in the form of Weights & Biases logs, and inference code. This first release includes four variants of our language model at the 7B scale corresponding to different architectures, optimizers, and training hardware, and one model at the 1B scale, all trained on at least 2T tokens. We are also releasing hundreds of intermediate checkpoints available as revisions on HuggingFace. For dataset building and analysis, it includes the full training data used for these models, including code that produces the training data, from AI2’s Dolma (Soldaini et al., 2024), and WIMBD (Elazar et al., 2023) for analyzing pretraining data. For evaluation, it includes AI2’s Catwalk (Groeneveld et al., 2023) for downstream evaluation and Paloma (Magnusson et al., 2023) for perplexity-based evaluation. For instruction-tuning, we released Open Instruct (Ivison et al., 2023; Wang et al., 2023), and we are currently using it to produce an adapted (instruction-tuned and RLHFed) version of OLMo, which we will release soon. Finally, all code and weights are released under the Apache 2.0 License. [^1] http://www.apache.org/licenses/LICENSE-2.0
This is the first step in a long series of planned releases, continuing with larger models, instruction-tuned models, and more modalities and variants down the line. We therefore hope to catalyze research into as-yet poorly understood aspects of these models, for example, the relationship between pretraining data and model capabilities, the impact of design and hyperparameter choices, and various optimization methods and their impact on model training. In addition, we report on the lessons learned and important details necessary to successfully train language models at this scale.
2 OLMo Framework
This section describes the OLMo framework, consisting of the OLMo models (Section 2.1), our pre-training dataset, Dolma (Section 2.2), and our evaluation framework (Section 2.4).
2.1 OLMo Model and Architecture
We adopt a decoder-only transformer architecture based on Vaswani et al. (2017), and deliver 1B and 7B variants as described in Table 1, with a 65B version coming soon. Our specific architecture includes several improvements over the vanilla transformer from Vaswani et al. (2017) following other recent large language models like PaLM (Chowdhery et al., 2022), the LLaMA family (Touvron et al., 2023a, b), OpenLM (Gururangan et al., 2023), and Falcon (Almazrouei et al., 2023). Table 2 gives a comprehensive comparison of our 7B architecture to the similarly-sized models from these other families.
Size Layers Hidden Size Attention Heads Tokens Trained
1B 16 2048 16 2T
7B 32 4086 32 2.46T
65B* 80 8192 64
Table 1: OLMo model sizes and the maximum number of tokens trained to.
                                                                                                                                    * At the time of writing our 65B model is still training.
We generally select hyperparameters by optimizing for training throughput on our hardware while minimizing the risk of loss spikes and slow divergence. We ablate choices through our in-loop evaluation setting, given available computational sources (Section 2.4). Table 2 compares our design choices with recent state-of-the-art open language models. Our main changes over the vanilla transformer architecture can be summarized as follows:
  1. 1.
    No biases. Following LLaMA, PaLM, and others, we exclude all bias terms from our architecture in order to improve training stability.
  2. 2.
    Non-parametric layer norm. We use the non-parametric formulation of layer norm (Ba et al., 2016) in which there is no affine transformation within the norm, i.e. no “adaptive gain” (or bias). We believe this was the safest option and it was also the fastest compared to the other variants we considered: parametric layer norm and RMSNorm (Zhang and Sennrich, 2019).
  3. 3.
    SwiGLU activation function. Like LLaMA, PaLM, and others we use the SwiGLU activation function (Shazeer, 2020) instead of ReLU, and following LLaMA the activation hidden size is approximately $\frac{8}{3}d$, but increased to the closest multiple of 128 (e.g. 11,008 for our 7B model) to improve throughput. [^2] Since SwiGLU is a “gated” activation function, the output is half the size of the input. So technically our inputs to SwiGLU have a dimensionality of 2 $\times$ 11,008 = 22,016 for our 7B model.
  4. 4.
    Rotary positional embeddings (RoPE). Like LLaMA, PaLM, and others we replace absolute positional embeddings with rotary positional embeddings (RoPE; Su et al., 2021).
  5. 5.
    Vocabulary. We use a modified version of the BPE-based tokenizer from GPT-NeoX-20B (Black et al., 2022) with additional tokens for masking personal identifiable information (PII). The final vocabulary size is 50,280. However, to maximize training throughput we increase the size of the corresponding embedding matrix in our model to 50,304 so that it’s a multiple of 128.
OLMo-7B LLaMA2-7B OpenLM-7B Falcon-7B PaLM-8B
Dimension 4096 4096 4096 4544 4096
Num heads 32 32 32 71 16
Num layers 32 32 32 32 32
MLP ratio $\sim$8/3 $\sim$8/3 $\sim$8/3 4 4
Layer norm type non-parametric RMSNorm parametric parametric parametric
Positional embeddings RoPE RoPE RoPE RoPE RoPE
Attention variant full GQA full MQA MQA
Biases none none in LN only in LN only none
Block type sequential sequential sequential parallel parallel
Activation SwiGLU SwiGLU SwiGLU GeLU SwiGLU
Sequence length 2048 4096 2048 2048 2048
Batch size (instances) 2160 1024 2048 2304 512
Batch size (tokens) $\sim$4M $\sim$4M $\sim$4M $\sim$4M $\sim$1M
Weight tying no no no no yes
Table 2: LM architecture comparison at the 7–8B scale. In the “layer norm type” row, “parametric” and “non-parametric” refer to the usual layer norm implementation with and without adaptive gain and bias, respectively.
2.2 Pretraining Data: Dolma
Despite progress in access to model parameters, pretraining datasets are still not as open. Pretraining data are often not released alongside open models (let alone closed models) and documentation about such data is often lacking in detail that would be needed to reproduce or fully understand the work. This has made it difficult to support certain threads of language model research, such as understanding how training data impacts model capabilities and limitations. To facilitate open research on language model pretraining, we built and released our pretraining dataset, Dolma—a diverse, multi-source corpus of 3T tokens across 5B documents acquired from 7 different data sources that are (1) commonly seen in large-scale language model pretraining and (2) accessible to the general public (Soldaini et al., 2024). Table 3 provides a high-level overview of the amount of data from each source.
Dolma is built using a pipeline of (1) language filtering, (2) quality filtering, (3) content filtering, (4) deduplication, (5) multi-source mixing, and (6) tokenization. We refer the reader to the Dolma report (Soldaini et al., 2024) for more details about its design principles, details about its construction, and a more detailed summary of its contents. The report provides additional analyses and experimental results from training language models on intermediate states of Dolma to share what we learned about important data curation practices, including the role of content or quality filters, deduplication, and mixing data from multiple sources. We keep documents from each source separate, both during curation as well as in the final release. We open-sourced our high-performance data curation tools; this toolkit can be used to further experiment on Dolma, reproduce our work, and enable fast and easy curation of pretraining corpora. Finally, we also open-sourced our WIMBD tool (Elazar et al., 2023) to help with dataset analysis.
Source Doc Type
UTF-8
bytes
Documents
GPT-NeoX
tokens
Common Crawl
web pages 9,022 3,370 2,006
The Stack
code 1,043 210 342
C4
web pages 790 364 174
Reddit
social media 339 377 80
peS2o
STEM papers 268 38.8 57
Project Gutenberg
books 20.4 0.056 5.2
Wikipedia, Wikibooks
encyclopedic 16.2 6.2 3.7
Total 11,519 4,367 2,668
Table 3: Composition of Dolma.
2.3 Adaptation
Pretrained models are not always used as-is, but rather further fine-tuned to improve their performance, safety, and usability. Often models are first trained to follow instructions (Mishra et al., 2022; Wei et al., 2022; Sanh et al., 2022), and then further trained on human preferences (Ouyang et al., 2022) to improve the quality of their generations. We showcase the efficacy of using OLMo as a base model for further fine-tuning by training OLMo to be a general chat assistant following our Open Instruct (Tülu) data and training setup (Ivison et al., 2023). Our approach involves first performing instruction fine-tuning with a mixture of distilled and human-written instruction data and then further aligning the model with distilled preference data using Direct Preference Optimization (DPO) (Rafailov et al., 2023). We experimented with mixing the Tulu instruction data at the end of pretraining, as done in recent models such as DeepSeek-AI et al. (2024), but did not have conclusive findings.
2.4 Evaluation
We perform base model evaluation at two stages: online evaluation to make decisions for model design and offline evaluation to evaluate model checkpoints. For the offline stage, we use the Catwalk framework (Groeneveld et al., 2023), a publicly available evaluation tool with access to a wide range of datasets and task formats. Using Catwalk, we perform downstream evaluation as well as intrinsic language modeling evaluation on the new perplexity benchmark, Paloma (Magnusson et al., 2023).
For both downstream and perplexity evaluation, we use our fixed evaluation pipeline to compare results against publicly available models. We also report a separate evaluation of our adapted model.
In-Loop Training Ablations
Throughout model training, we perform downstream evaluations to make decisions around model architecture, initialization, optimizers, learning rate schedule, and data mixtures. We call this our online evaluation as it runs in-loop every 1000 training steps (or $\sim$4B training tokens) and provides an early and continuous signal on the quality of the model being trained. These evaluations rely on many of the core tasks and experiment settings used for our offline evaluation detailed in Section 4.1, which also mirrors the task and evaluation structure of the EleutherAI eval harness (Gao et al., 2023).
Downstream Evaluation
Following much previous work (Brown et al., 2020; Black et al., 2022; Touvron et al., 2023a, b, inter alia), we report zero-shot performance on a set of downstream tasks. Our evaluation suite consists of 8 core tasks corresponding closely to the commonsense reasoning task set reported by Touvron et al. (2023a) and Touvron et al. (2023b) (see Table 6 for a list of tasks). Given the scale of the models being evaluated, such tasks were selected at the beginning of model development due to their naturalness (e.g., all can formulated as text completion scoring tasks) and ability to provide meaningful signals throughout training (see Figure 1).
Intrinsic Language Modeling Evaluation
To measure how OLMo-7B fits distributions of language beyond held-out training data, we use Paloma (Magnusson et al., 2023), a new perplexity benchmark that includes 585 different domains of text. Domains range from nytimes.com to r/depression on Reddit and are drawn from 18 separate data sources, such as C4 (Raffel et al., 2020), in stratified samples. This allows for more equal inclusion of text domains that are under-represented in their source corpora.
We aim not just to compare OLMo-7B against other models for best performance, but also to demonstrate how it enables fuller and more controlled scientific evaluations. OLMo-7B is the largest LM with explicit decontamination for perplexity evaluation. Following the approach described in Paloma, we remove any pretraining document with paragraphs leaked from Paloma evaluation data. Without decontamination, other models risk underestimating perplexity (i.e., overestimating the model’s out-of-sample fit). We also release intermediate checkpoints, allowing richer comparisons with two other models that release checkpoints, Pythia-6.9B (Biderman et al., 2023) and RPJ-INCITE-7B (Together Computer, 2023) (see Figure 2).
Adaptation Evaluation
We also follow our Open Instruct evaluation suite Wang et al. (2023); Ivison et al. (2023) to evaluate OLMo after instruction fine-tuning and DPO training using our We focus on evaluations around model chat capabilities and safety to showcase the efficacy of using OLMo as a base for further fine-tuning.
3 Training OLMo
This section describes our pretraining setup, including our distributed training framework (Section 3.1), optimizer settings (Section 3.2), data preparation (Section 3.3), and hardware (Section 3.4).
3.1 Distributed Training Framework
We train our models using the ZeRO optimizer strategy (Rajbhandari et al., 2019) via PyTorch’s FSDP framework (Zhao et al., 2023), which reduces memory consumption by sharding the model weights and their corresponding optimizer state across GPUs. At the 7B scale, this enables training with a micro-batch size of 4096 tokens per GPU on our hardware (see Section 3.4). For OLMo-1B and -7B models, we use a constant global batch size of approximately 4M tokens (2048 instances, each with a sequence length of 2048 tokens). For OLMo-65B model (currently training), we use a batch size warmup that starts at approximately 2M tokens (1024 instances), then doubles every 100B tokens until reaching approximately 16M tokens (8192 instances).
To improve throughput, we employ mixed-precision training (Micikevicius et al., 2017) through FSDP’s built-in settings and PyTorch’s amp module. The latter ensures that certain operations like the softmax always run in full precision to improve stability, while all other operations run in half-precision with the bfloat16 format. Under our specific settings, the sharded model weights and optimizer state local to each GPU are kept in full precision. The weights within each transformer block are only cast to bfloat16 when the full-sized parameters are materialized on each GPU during the forward and backward passes. Gradients are reduced across GPUs in full precision.
3.2 Optimizer
We use the AdamW optimizer (Loshchilov and Hutter, 2019) with the hyperparameters shown in Table 4. For all model sizes, we warm up the learning rate over 5000 steps ($\sim$21B tokens) and then decay it linearly from there down to a tenth of the peak learning rate over the remainder of training. After the warm-up period, we clip gradients such that the total $l^{2}$-norm of the parameter gradients [^3] During gradient clipping all of the model’s parameters are treated as a single big vector (as if all parameters were flattened and concatenated together), and we take the $\ell_{2}$-norm over the corresponding single gradient vector. This is the standard way to clip gradients in PyTorch. does not exceed $1.0$. Table 5 gives a comparison of our optimizer settings at the 7B scale to those of other recent LMs that also used AdamW.
3.3 Data
We built our training dataset out of a 2T-token sample from our open dataset, Dolma (Soldaini et al., 2024), which we describe in Section 2.2. The tokens from every document are concatenated together after appending a special EOS token to the end of each document, and then we group consecutive chunks of 2048 tokens to form training instances. The training instances are shuffled in the exact same way for each training run. The data order and exact composition of each training batch can be reconstructed from the artifacts we release.
All of our released models have been trained to at least 2T tokens (a single epoch over our training data), and some have been trained beyond that by starting a second epoch over the data with a different shuffling order. The impact of repeating this small amount of data should be negligible according to prior work (Muennighoff et al., 2023).
Size Peak LR Betas Epsilon Weight Decay Batch Size (tokens)
1B 4.0E-4 (0.9, 0.95) 1.0E-5 0.1 $\sim$4M
7B 3.0E-4 (0.9, 0.95) 1.0E-5 0.1 $\sim$4M
65B* 1.5E-4 (0.9, 0.95) 1.0E-5 0.1 $\sim$2M $\rightarrow$ $\sim$4M $\rightarrow$ $\sim$8M $\rightarrow$ $\sim$16M
Table 4: AdamW pretraining hyperparameters for OLMo models.
                                                                                                                                    * At the time of writing our 65B model is still training.
OLMo-7B LLaMA2-7B OpenLM-7B Falcon-7B
warmup steps 5000 2000 2000 1000
peak LR 3.0E-04 3.0E-04 3.0E-04 6.0E-04
minimum LR 3.0E-05 3.0E-05 3.0E-05 1.2E-05
weight decay 0.1 0.1 0.1 0.1
beta1 0.9 0.9 0.9 0.99
beta2 0.95 0.95 0.95 0.999
epsilon 1.0E-05 1.0E-05 1.0E-05 1.0E-05
LR schedule linear cosine cosine cosine
gradient clipping global 1.0 global 1.0 global 1.0 global 1.0
gradient reduce dtype FP32 FP32 FP32 BF16
optimizer state dtype FP32 most likely FP32 FP32 FP32
Table 5: Comparison of pretraining optimizer settings at the 7B scale. Each model in this table used AdamW as its optimizer.
3.4 Hardware
In order to verify that our codebase could be used on both NVIDIA and AMD GPUs without any loss in performance, we trained models on two different clusters:
  • LUMI: Provided by the LUMI supercomputer, [^4] https://www.lumi-supercomputer.eu we used up to 256 nodes on this cluster, where each node consists of 4x AMD MI250X GPUs with 128GB of memory [^5] The MI250X is a dual-chip module, meaning in practice that each physical device consists of two logical devices, so each node has 8 logical GPU devices with 64GB of memory each. and 800Gbps of interconnect.
  • MosaicML: Provided by MosaicML [^6] https://www.mosaicml.com (Databricks), we used 27 nodes on this cluster, where each node consists of 8x NVIDIA A100 GPUs with 40GB of memory and 800Gbps interconnect.
Despite minor differences in batch size to optimize for training throughput, both runs resulted in nearly identical performance on our evaluation suite by 2T tokens.
4 Results
The checkpoint used for evaluating OLMo-7B is trained until 2.46T tokens on the Dolma (Soldaini et al., 2024) dataset with a linear learning rate decay schedule mentioned in Section 3.2. In our experiments, we find that tuning this checkpoint further on the Dolma dataset for 1000 steps with the learning rate linearly decayed to 0 boosts model performance on perplexity and end-task evaluation suites described in Section 2.4. We compare OLMo with other publicly available models including LLaMA-7B (Touvron et al., 2023a), LLaMA2-7B (Touvron et al., 2023b), MPT-7B (MosaicML NLP Team, 2023), Pythia-6.9B (Biderman et al., 2023), Falcon-7B (Almazrouei et al., 2023) and RPJ-INCITE-7B (Together Computer, 2023).
4.1 Downstream evaluation
Setup
Our core downstream evaluation suite (see Table 6) consists of: arc (both arc_easy and arc_challenge) (Clark et al., 2018), boolq (Clark et al., 2019), openbookqa (Mihaylov et al., 2018), sciq (Welbl et al., 2017), hellaswag (Zellers et al., 2019), piqa (Bisk et al., 2020), and winogrande (Sakaguchi et al., 2021). In Appendix A, we also report results on an additional set of auxiliary tasks outside of our core evaluation set that we found to have less stable performance trends (see Figure 4).
In all cases, we perform zero-shot evaluation using the rank classification approach popularized by Brown et al. (2020). Under this approach, candidate text completions (e.g., different multiple-choice options) are ranked by likelihood (usually normalized by some normalization factor), and prediction accuracy is reported. While Catwalk implements several common likelihood normalization strategies, including normalizing by number of tokens (per-token normalization) (Brown et al., 2020; Liang et al., 2022), by number of characters (per-character normalization) (Gao et al., 2023), as well as incorporating an answer’s unconditional likelihood (Brown et al., 2020), we selected the normalization strategies for each dataset separately. Specifically, we used unconditional normalization for arc and openbookqa, per-token normalization for hellaswag, piqa, and winogrande and no normalization for boolq, and sciq (i.e., tasks formulated as single token prediction tasks).
7B Models
arc
challenge
arc
easy
boolq
hella-
swag
open
bookqa
piqa sciq
wino-
grande
avg.
Falcon 47.5 70.4 74.6 75.9 53.0 78.5 93.9 68.9 70.3
LLaMA 44.5 67.9 75.4 76.2 51.2 77.2 93.9 70.5 69.6
Llama 2 48.5 69.5 80.2 76.8 48.4 76.7 94.5 69.4 70.5
MPT 46.5 70.5 74.2 77.6 48.6 77.3 93.7 69.9 69.8
Pythia 44.1 61.9 61.1 63.8 45.0 75.1 91.1 62.0 63.0
RPJ-INCITE 42.8 68.4 68.6 70.3 49.4 76.0 92.9 64.7 66.6
[] OLMo-7B 48.5 65.4 73.4 76.4 50.4 78.4 93.8 67.9 69.3
Table 6: Zero-shot evaluation of OLMo-7B and 6 other publicly available comparable model checkpoints on 8 core tasks from the downstream evaluation suite described in Section 2.4. For OLMo-7B, we report results for the 2.46T token checkpoint.
Results
Table 6 summarizes the result of zero-shot evaluation of OLMo-7B and compares it against 6 other publicly available models of comparable size. We report results on 8 core tasks from our evaluation suite described in Section 2.4. On aggregate, OLMo-7B is competitive against all 6 publicly available model checkpoints in our comparison table.
In Figure 1 we plot the accuracy score progression of 8 core end-tasks. All tasks, except OBQA, show an upward trend in accuracy numbers as OLMo-7B is trained on more tokens. A sharp upward tick in accuracy of many tasks between the last and the second to last step shows us the benefit of linearly reducing the LR to 0 over the final 1000 training steps. See Table 9 in Appendix A for additional evaluation results and discussion.
Figure 1: Accuracy score progression of OLMo-7B on 8 core end-tasks score from Catwalk evaluation suite described in Section 2.4. We can see the benefit of decaying LR to 0 in the final 1000 steps of training on most tasks.
4.2 Intrinsic language modeling evaluation
Setup
For intrinsic evaluations, Paloma proposes a range of analyses, from inspection of performance in each domain separately to more summarized results over combinations of domains. We report results at two levels of granularity: the aggregate performance over 11 of the 18 sources in Paloma as in Magnusson et al. (2023), as well as more fine-grained results over each of these sources individually. This particular subset of 11 sources from Paloma excludes sources that are not publicly available, involve fringe or toxic text, or consist of code data not supported by Paloma’s decontamination approach. This leaves C4 (Raffel et al., 2020), mC4-en (Chung et al., 2023), Wikitext 103 (Merity et al., 2016), Penn Treebank (Marcus et al., 1999; Nunes, 2020), RedPajama (Together Computer, 2023), Falcon-RefinedWeb (Penedo et al., 2023), Dolma (Soldaini et al., 2024), M2D2 S2ORC (Reid et al., 2022), M2D2 Wikipedia (Reid et al., 2022), C4 100 domains (Chronopoulou et al., 2022), and Dolma 100 Subreddits (Soldaini et al., 2024). To allow for a fair comparison between models with different vocabularies, we report bits per byte as defined by Gao et al. (2020) over the test sets of these sources.
Results
In the Sources Combined subplot of Figure 2, we show the performance of OLMo-7B against 6 comparably-sized language models on the combination of 11 data sources from Paloma. Overall we find OLMo to have a competitive fit, especially given its training data was explicitly decontaminated against Paloma. As seen through the comparison of final models (see shapes) as well intermediate checkpoints (see dashed lines), the OLMo results follow similar scaling trends of other models. Note that the performance of intermediate checkpoints is influenced by where that checkpoint occurs in the learning rate schedule. So models trained for fewer steps will tend to have steeper training curves without necessarily being more sample efficient if training duration were fixed across all models. MPT-7B, nevertheless, stands out as improving ahead of the other models in this subplot. This could be due to a number of factors, including pretraining data composition and its match to the domains in Paloma (e.g., MPT trains on 27% non-Common Crawl data rather than 18% for LLaMA, 12.2% for RedPajama, and 11.2% for OLMo) as well as various data preprocessing decisions (e.g., MPT’s use of semantic deduplication by Abbas et al., 2023, on C4).
Figure 2: Bits per byte on 11 evaluation data sources from Paloma and their combination (Magnusson et al., 2023), decontaminated from OLMo’s pretraining data. While models follow a general data scaling trend, sample efficiency is most favorable on in-distribution data. For example, OLMo-7B overtakes all other models on C4, perhaps from having 88.8% Common Crawl pretraining data.
The remaining subplots in Figure 2 provide more fine-grained analysis by reporting bits per byte separately for each of the 11 data sources that are combined in the aggregated Paloma metric. From this we see greater variation in sample efficiency, largely driven by the similarity of training and evaluation distributions. Notably, OLMo-7B fares well on evaluations predominated by Common Crawl, such as C4, though different ways of postprocessing Common Crawl are best fit by models trained with that specific data, such as Falcon-7B on Falcon RefinedWeb. Meanwhile, OLMo-7B is less sample efficient compared to other models on sources less related to scraped web text, such as WikiText-103, M2D2 S2ORC, and M2D2 Wikipedia. The RedPajama evaluation shows a similar pattern, perhaps as only 2 of its 7 domains are from Common Crawl, and Paloma weights domains within each source equally. Since heterogeneous data from curated sources like Wikipedia and ArXiv papers is much less abundant than scraped web text, maintaining sample efficiency for fit to these distributions of language will be challenging as pretraining corpora are scaled.
4.3 Adaptation Evaluation
Model MMLU AlpacaEval ToxiGen TruthfulQA 0-shot $\uparrow$ %win $\uparrow$ % Toxic $\downarrow$ %Info+True $\uparrow$ OLMo (base) 28.3 - 81.4 31.6 MPT Chat 33.8 46.8 0.1 42.7 Falcon Instruct 25.2 14.0 70.7 27.2 RPJ-INCITE Chat 27.0 38.0 46.4 53.0 Llama-2-Chat 46.8 87.3 0.0 26.3 Tülu 2 50.4 73.9 7.0 51.7 Tülu 2+DPO 50.7 85.1 0.5 - * [] OLMo +SFT 47.3 57.0 14.4 41.2 [] OLMo +SFT+DPO 46.2 69.3 1.7 52.0
Table 7: Evaluation of various instruction-tuned 7B models, including OLMo-7B and before and after adaptation training. Lower is better for ToxiGen and higher is better for other metrics. We provide a detailed description of models and metrics in Appendix. C. * Following Ivison et al. (2023), we do not report Tülu 2 TruthfulQA scores due to test set contamination.
Setup
We evaluate OLMo before adaptation, and after both the supervised fine-tuning and DPO training stage, focusing on the safety and chat evaluations used by Wang et al. (2023). We additionally compare to officially released instruction-tuned variants of the models from Table 6. We finally also compare to Tülu 2 models to compare against models trained using the same post-training data mixes and procedures.
Results
We find that instruction tuning considerably improves the performance and safety of OLMo, increasing MMLU performance by a wide margin and improving ToxiGen and TruthfulQA scores - especially after DPO training. Additionally, we find that OLMo outperforms most other chat variants after both initial instruction tuning (OLMo +SFT) and additional preference alignment (OLMo +SFT+DPO), highlighting both the strength of OLMo as a base model and the strength of the Tülu mix used to perform adaptation training. However, we find there is still a gap with Tülu 2, which is trained by applying the Tülu mix on Llama 2. This gap may be due to test set contamination in Llama 2777Touvron et al. (2023b) report that Llama 2 was pretrained on data contaminated with MMLU test data. and because the Tülu mix was primarily designed for Llama models - we will investigate the cause of this gap in future work. Overall, we see that OLMo greatly benefits from additional tuning and serves as a strong base model for downstream applications.
4.4 Power Consumption and Carbon Footprint
Following previous literature (Strubell et al., 2019; Patterson et al., 2021; Wu et al., 2022; Dodge et al., 2022), we estimate the total energy consumed and carbon released while pretraining our models by calculating the total power consumption required for training, and then multiplying it by the carbon emission intensity of the power grid where the model was trained. While reporting these operational emissions is standard practice, it does not account for other sources of emissions such as the embodied emissions due to the manufacturing, transportation and disposal of hardware and datacenter infrastructure, lifetime operational emissions due to use, rebound effects, or other environmental impacts such as water consumption or mining. Thus our estimates should be viewed as lower bounds.
We calculate the total power consumption for our models by measuring the power consumption of a single node every 25ms, calculating an average across the entire training run, and multiplying by the total number of nodes. We then account for the energy efficiency of the data center by multiplying the previous total by a power usage effectiveness (PUE) factor, which we set to 1.1, representing a conservative 10% energy consumption overhead typical of energy efficient datacenters. [^8] https://www.nrel.gov/computational-science/measuring-efficiency-pue.html [^9] https://www.google.com/about/datacenters/efficiency/ We estimate that pretraining our 7B models consumed 239 MWh of energy.
[…]
We hope that openly releasing our models can reduce future emissions by allowing others to avoid the need to pretrain models from scratch, and give insights into the true cost of developing state of the art models. We also highlight that our estimates are lower bounds, because they do not include other critical pieces of development such as debugging, hyperparameter tuning, and downtime.
GPU Type
GPU Power
Consumption
Power
Usage
Effectiveness
Carbon
Intensity
[…]
Carbon
Emissions
[…]
Gopher-280B TPU v3 1,066 1.08 0.330 380
BLOOM-176B A100-80GB 433 1.2 0.057 30
OPT-175B A100-80GB 324 1.1 0.231 82
T5-11B TPU v3 77 1.12 0.545 47
LLaMA-7B A100-80GB 33 1.1 0.385 14
LLaMA2-7B A100-80GB 74 1.1 0.385 31
[] OLMo-7B MI250X 135 1.1 $0$*
[] OLMo-7B A100-40GB 104 1.1 0.610 70
[…]
5 Artifacts Released
By sharing artifacts from all pipeline stages, we aim to encourage open research and reduce duplicated, often costly efforts, by academics and practitioners. We release the following:
  1. 1.
    The training and modeling code. [^13] https://github.com/allenai/OLMo
  2. 2.
    The trained model weights for the 7B model, [^14] https://huggingface.co/allenai/OLMo-7B 7B-twin-2T, [^15] https://huggingface.co/allenai/OLMo-7B-Twin-2T and the 1B model. [^16] https://huggingface.co/allenai/OLMo-1B For all the models, we release not only the final model weights but also 500+ intermediate checkpoints at intervals of 1000 steps.
  3. 3.
    Adapted OLMo-7B with instruction-tuning, 7B-SFT [^17] https://huggingface.co/allenai/OLMo-7B-SFT, and RLHF, 7B-Instruct [^18] https://huggingface.co/allenai/OLMo-7B-Instruct including its training and evaluation code and data using our Open Instruct [^19] https://github.com/allenai/open-instruct library (Wang et al., 2023; Ivison et al., 2023).
  4. 4.
    The training data Dolma (Soldaini et al., 2024). [^20] https://huggingface.co/datasets/allenai/dolma
  5. 5.
    Dolma’s toolkit to construct new datasets, [^21] https://github.com/allenai/dolma and WIMBD (Elazar et al., 2023) for dataset analysis. [^22] https://github.com/allenai/wimbd
  6. 6.
    The evaluation code [^23] https://github.com/allenai/OLMo-Eval using Catwalk [^24] https://github.com/allenai/catwalk for downstream evaluation (Groeneveld et al., 2023) and Paloma [^25] https://paloma.allen.ai for perplexity-based evaluation (Magnusson et al., 2023).
  7. 7.
    The complete set of metrics logged to Weights & Biases during training. [^26] https://wandb.ai/ai2-llm/OLMo-7B/reports/OLMo-7B--Vmlldzo2NzQyMzk5
We intend to follow up on this release with further training logs, ablations, and findings.
6 License
Our goal is to facilitate scientific development and empower the scientific community, so we favor permissive licenses that give users flexibility in using our resources and artifacts. As such, all code and weights are released under the Apache 2.0 License. [^27] http://www.apache.org/licenses/LICENSE-2.0 Some licenses used by other organizations for recent model releases prohibit using the outputs from their models to train artificial intelligence or machine learning systems, while we expressly allow users to do so. We also do not limit commercial use. We hope that our models can make other models better. We recognize that the risk for misuse of our models is relatively low since they are mainly designed as scientific artifacts not as products with broad public adoption (our models have not been adapted as chatbots). In addition, over the past year there have been a number of comparable models released with very permissive licenses, so using a more strict license for our work will not remove the overall risk in the field. We believe this tradeoff on the side of being more open is the best option.
7 Conclusion and Future Work
This technical report presents our first release of OLMo, a state-of-the-art, truly open language model and its framework to build and study the science of language modeling. Unlike most prior efforts that have only released model weights and inference code, we release OLMo and the whole framework, including training data and training and evaluation code. Soon, we will also release training logs, ablations, findings and Weights & Biases logs. We are also exploring the adaptation of OLMo with instruction tuning and different flavors of RLHF. We are going to release the adapted models as well as all of our model adaptation code and data.
We intend to continuously support and extend OLMo and its framework, and continue to push the boundaries of open LMs to empower the open research community. To that end, we look forward to bringing different model sizes, modalities, datasets, safety measures, and evaluations into the OLMo family. We hope this and future releases will empower and strengthen the open research community and inspire a new wave of innovation.
Author Contributions
OLMo would not have been possible without the help of our many teammates and collaborators. We list author contributions (in alphabetical order) below:
Contributors to pretraining dataset construction and tooling (Dolma) include Russell Authur, Iz Beltagy, Akshita Bhagia, Khyathi Chandu, Jesse Dodge, Yanai Elazar, Dirk Groeneveld, Rodney Kinney, Kyle Lo, Aakanksha Naik, Abhilasha Ravichander, Dustin Schwenk, Luca Soldaini, and Nishant Subramani.
Contributors to model training and architecture include Shane Arora, Iz Beltagy, Akshita Bhagia, Matthew E. Peters, Dirk Groeneveld, Ananya Harsh Jha, William Merrill, Jacob Morrison, Niklas Muennighoff, Dustin Schwenk, Saurabh Shah, Pete Walsh, and Mitchell Wortsman.
Contributors to evaluation suite and tooling include Akshita Bhagia, Arman Cohan, Pradeep Dasigi, Jesse Dodge, Dirk Groeneveld, Yuling Gu, Tushar Khot, Ian Magnusson, Kyle Richardson, Oyvind Tajford, and Pete Walsh.
Contributors to model adaptation include Iz Beltagy, Pradeep Dasigi, Jack Hessel, Hamish Ivison, Nathan Lambert, Valentina Pyatkin, Pete Walsh, and Yizhong Wang.
Contributors to license creation and risk assessment include David Atkinson, Jesse Dodge, Jennifer Dumas, Crystal Nam, and Will Smith.
The OLMo project was led by Hannaneh Hajishirzi and Noah A. Smith.
Acknowledgements
OLMo would not have been possible without the support of many individuals and institutions. The experimental components of this work were made possible through a partnership with AMD and CSC, enabling use of the LUMI supercomputer, and Kempner Institute at Harvard University. We thank Jonathan Frankle and the team at MosaicML (now Databricks) for sharing their experiences with FSDP, and building the code base that OLMo is based on. We thank our teammates Taira Anderson, Michelle Benedict, Jon Borchardt, Evie Cheng, Arnavi Chheda, Johann Dahm, Matt Latzke, Kelsey MacMillan, Aaron Sarnat, Carissa Schoenick, Sam Skjonsberg, Michael Schmitz, Michael Wilson, Caitlin Wittlif, and the entire IT team, for their help with the website, design, internal and external communications, budgeting, and other activities that supported smooth progress on this project. Finally, we also express gratitude for the helpful discussions and feedback from our teammates at AI2 and close collaborators, including Prithviraj (Raj) Ammanabrolu, Peter Clark, Nicole DeCario, Doug Downey, Ali Farhadi, Ian Ferreira, Väinö Hatanpää, Sham M. Kakade, Julien Launay, Sydney Levine, Pekka Manninen, Franzi Roessner, Maarten Sap, Ludwig Schmidt, Yulia Tsvetkov, and Daniel S. Weld.
Appendix A Additional Evaluation
Figure 3: Bits per byte for each of the 7 remaining Paloma data sources not aggregated in Figure 2.
Additional perplexity results
In Figure 3 we provide results for each of the 7 data sources in Paloma [Magnusson et al., 2023] that are excluded from the combined metric in Figure 2. Some of these sources such as Pile [Gao et al., 2020] and ICE [Greenbaum and Nelson, 1996] are not publicly available at this time. Dolma 100 Programming Languages [Soldaini et al., 2024] consists of code data that is not supported by the decontamination approach used in Paloma. TwitterAAE [Blodgett et al., 2016], along with ICE, are datasets for targeted analyses of disparities in performance between different dialects and as such should be evaluated separately. And finally, the Manosphere, Gab, and 4chan corpora [Ribeiro et al., 2021, Zannettou et al., 2018, Papasavva et al., 2020] are intended to examine model fit to language from fringe online communities that are studied for prevalent hate speech and toxicity. Thus minimizing perplexity on these fringe corpora is not always desirable.
One notable result here is that OLMo-7B is much farther ahead of the other models on Dolma 100 Programming Languages (100 PLs). Note that this effect may be due in part to underestimation from contamination, as decontaminating code data is beyond the scope of the method in Paloma. At the same time other models that are trained on code data from GitHub such as RPJ-INCITE-7B, that are just as likely to have contamination, fair much worse. Another factor then is that OLMo-7B trains on code data with exactly the same post-processing as that in 100 PLs while the code data in other models will have been processed differently. Similarly, Pile evaluation demonstrates these in-distribution and potential contamination effects as Pythia-6.9B achieves top performance despite being trained on almost an order of magnitude fewer tokens than OLMo-7B.
The results on the remaining 5 targeted sources should be interpreted with care, as Paloma often finds that perplexity on these sources is dominated by superficial features such as low average document length rather than fit to that which would actually be salient to members of these speech communities. TwitterAAE and Gab have among the shortest documents in Paloma contributing to unusually high bits per byte in this figure. Other than these two, the models are notably very closely grouped in a data scaling trend in ICE, Manosphere, and 4chan.
Additional end-task results
Next, in Table 9, we provide results from zero-shot evaluation of OLMo-7B on 6 additional end-tasks apart from the 8 in our core evaluation suite. These tasks are headqa_en [Vilares and Gómez-Rodríguez, 2019], logiqa [Liu et al., 2020], mrpc [Dolan and Brockett, 2005], qnli [Wang et al., 2018], wic [Pilehvar and Camacho-Collados, 2018], and wnli [Wang et al., 2018].
headqa_en logiqa mrpc qnli wic wnli avg. Falcon-7B 38.6 23.7 62.8 49.8 49.5 47.9 45.4 LLaMA-7B 38.7 19.5 68.6 50.1 49.1 52.1 46.4 LLaMA2-7B 39.5 26.1 69.1 49.4 49.8 45.1 46.5 MPT-7B 37.4 22.9 67.7 52.1 48.1 47.9 46.0 Pythia-6.9B 40.1 21.5 65.4 53.8 55.0 38.0 45.6 RPJ-INCITE-7B 36.9 27.8 58.8 53.8 48.9 57.8 47.3 [] OLMo-7B 37.3 23.4 68.4 49.1 50.2 56.3 47.5
Table 9: Zero-shot evaluation of OLMo-7B on 6 additional end-tasks apart from the 8 present in our core evaluation suite. Once again, we compare OLMo-7B to 6 other model checkpoints which are publicly available. We find that OLMo-7B outperforms the other models on aggregate taken over 6 additional end-tasks from this table, however these tasks were also found to provide limited signal during training (see Figure 4).
We note, however, that in contrast to our core evaluation set described in Section 4.1, we found these additional end-tasks to have less stable performance during model development, and to provide a limited signal. This is illustrated in Figure 4, where we see the progress of task performance throughout training to be more random (compare with the more stable upward trends in Figure 1). While tasks such as mrpc and wic appear more stable, they offered additional difficulties related to performance being tied to random chance (e.g., wic) or the tendency of models to make spurious predictions (e.g., always predicting a single label) that either inflate or deflate performance due to dataset class imbalances (e.g., mrpc). We therefore caution against relying too heavily on these tasks when measuring model performance throughout training and comparing models.
Figure 4: Accuracy score progression of OLMo-7B on 6 additional end-tasks. The performance of these additional end-tasks was unstable and provided limited signal during model development.
Appendix B Adaptation Training Details
We use the following hyperparameters when instruction tuning OLMo. These were chosen through small pilot experiments.
  • Learning Rate: $2\times 10^{-6}$
  • Epochs: 3
  • Warmup: Linear warmup for the first 3% of total training time, and then linear cooldown to a learning rate of 0 over the remaining steps.
  • Weight Decay: 0
  • Gradient clipping: 0
  • Maximum sequence length: 2048
After instruction finetuning, we then use the following hyperparameters for DPO training, following Ivison et al. [2023]:
  • Learning Rate: $5\times 10^{-7}$
  • $\beta$: 0.1
  • Epochs: 3
  • Warmup: Linear warmup for the first 10% of total training time, and then linear cooldown to a learning rate of 0 over the remaining steps.
  • Weight Decay: 0
  • Gradient clipping: 0
  • Maximum sequence length: 2048
Appendix C Adaptation Evaluation and Model details
We choose the models in Table 7 by choosing the ‘canonical’ best versions (that is, the best instruction-tuned or otherwise adapted models released by the same organisation) of the base models we compare against in Table 6. We additionally compare to Tülu 2 to show the current best models trained using the Tülu mix used to finetune OLMo. We display evaluations on MMLU, AlpacaEval, ToxiGen, and Truthfulness to focus on displaying how instruction tuning can generally help capabilities (MMLU), how the models perform in an open-ended chat setting (AlpacaEval), and to test how instruction tuning aids in model safety and truthfulness (AlpacaEval, ToxiGen). We additionally report OLMo’s performance over the entire Tülu evaluation suite in Table 10.
Model MMLU GSM8k BBH TydiQA Codex-Eval AlpacaEval ToxiGen TruthfulQA 0-shot 8-shot CoT 3-shot CoT 1-shot Pass@10 %win % Toxic % Info + True OLMo-7B 28.3 8.5 31.7 32.3 21.4 - 81.4 31.6 +SFT 47.3 15.5 36.9 35.2 28.6 57.0 14.4 41.2 +SFT+DPO 46.1 11.0 35.8 21.7 27.8 69.3 1.7 52.0
Table 10: Evaluation of OLMo-7B models before and after instruction finetuning and DPO training on the full Tülu evaluation suite. Lower is better for ToxiGen and higher is better for other metrics.
We provide a brief description of each model evaluated in Table 7 below. For all models, we use the provided chat template for prompt formatting when available.
  • MPT Chat: A version of MPT 7B finetuned on the ShareGPT-Vicuna [Chiang et al., 2023], HC3 [Guo et al., 2023], Alpaca [Taori et al., 2023], HH-RLHF [Bai et al., 2022], and Evol-Instruct [Xu et al., 2024] datasets. Retrieved from https://huggingface.co/mosaicml/mpt-7b-chat.
  • Falcon Instruct: A version of Falcon 7B finetuned on the Baize [Xu et al., 2023], GPT4All [Anand et al., 2023], GPTeacher [Teknium1, 2023], and Refined-Web English [Penedo et al., 2023] datasets. Retrieved from https://huggingface.co/tiiuae/falcon-7b-instruct.
  • RPJ-INCITE Chat: A version of RPJ-INCITE 7B finetuned on the OASST1 [Köpf et al., 2023] and Dolly V2 [Conover et al., 2023] datasets. Retrieved from https://huggingface.co/togethercomputer/RedPajama-INCITE-7B-Chat.
  • Llama-2 Chat: A version of Llama 2 7B finetuned on a mixture of instruction datasets and further trained with RLHF. We refer the reader to Touvron et al. [2023b] for further details.
  • Tülu 2: A version of Llama 2 7B finetuned on a mixture of instruction datasets (the Tülu 2 mix). We refer the reader to Ivison et al. [2023] for further details.
  • Tülu 2+DPO: Tülu 2 further trained with DPO on the UltraFeedback dataset [Cui et al., 2023]. We refer the reader to Ivison et al. [2023] for further details.
  • OLMo +SFT: A version of OLMo 7B fintuned on the same data as Tülu 2.
  • OLMo +SFT+DPO: OLMo +SFT further trained with DPO on the UltraFeedback dataset [Cui et al., 2023].
We additionally provide a brief description of each evaluation setting from Table 7:
  • MMLU: We use the official MMLU [Hendrycks et al., 2021] evaluation script and prompts available at https://github.com/hendrycks/test, with modifications to allow for batch processing. We evaluate using 0 few-shot examples, following the original setup of MMLU. We report average accuracy across test examples.
  • ToxiGen: We follow the setup in Touvron et al. [2023b], but use the original set of prompts from Hartvigsen et al. [2022], which are designed to elicit toxic generations for certain groups. We take only the prompts designed to produce toxic language (‘hateful’ prompts) and use 500 prompts per group to reduce evaluation costs. For base language models, we pass in the original ToxiGen prompts unchanged and greedily decode up to the first new line (or a maximum of 512 tokens). For instruction-tuned models, we place the prompt in the corresponding template, and ask the model to complete the prompt, until the model generates a stop token (or a maximum of 512 tokens). We pass the generated text into a roberta-large model trained to detect toxic content finetuned as part of Hartvigsen et al. [2022] [^28] https://huggingface.co/tomh/toxigen_roberta. We then report the percentage of generations deemed toxic by the classifier.
  • TruthfulQA: Following Touvron et al. [2023b], we mainly use the generation setting of TruthfulQA [Lin et al., 2022]. The TruthfulQA dataset contains 818 questions, which are used to prompt the tested model to generate answers. We use the default QA prompt format with 6 in-context QA examples. We follow the official script in their official implemention [^29] https://github.com/sylinrl/TruthfulQA/ to do greedy decoding and answer postprocessing. We train two LLaMA 2-based classifiers for judging the truthfulness and informativeness of the model response, due to the deprecation of GPT-3 making exact replication of the original TruthfulQA evaluation infeasible. We find that the LLaMA 2 judges are generally able to match the performance of the original GPT-3-based judges used by Lin et al. [2022]. We report the rate of the responses being truthful and informative (Informative and Truthful) following Touvron et al. [2023b]. We only report the % Informative and Truthful as our primary metric.
  • AlpacaEval: We use the package provided by Li et al. [2023], following the default setup which asks the evaluated model to generate responses for 805 prompts and employ GPT-4 to compare the response with Davinci-003. We employ the “alpaca_eval_gpt4” annotator. We allow the evaluated model to generate up to 2048 tokens, without specifying special stop sequences. The reported win-rate is the percentage of model generations that GPT-4 reports as being preferred over the generations from Davinci-003.